Emma paid the $10.02 bill for her lunch with 234 coins consisting of pennies, nickels and dimes. If the number of nickels plus dimes was equal to the number of pennis then how many coins of each type did she use?
She used ___pennis, ____, nickels, _____,dimes

Answer by ikleyn(52797)

.

From the condition, the number of pennies is half of 234, i.e.  234/2 = 117.


Thus the number of nickels PLUS dimes is 117, too, and these two type of coins are worth  1002 - 117 = 885 cents.


Now, if the number of dimes is x, then the number of nickels is  (117-x),

and the money equation is


    5*(117-x) + 10x = 885  cents,    or

    
     117-x + 2x = 885/5 = 177,

     x = 177 - 117 = 60.


ANSWER.  117 pennies, 60 dimes  and  117-60 = 57 nickels.


CHECK.   a)  117 + 60 + 57 = 234  coins.           ! Correct !

         b)  117 + 60*10 + 5*57 = 1002  cents.     ! Precisely correct !

Solved.

----------------

This problem seems to be on 3 equations in 3 unknowns.

But using the logic and a technique, presented in my post, it admits reducing to one single equation in one unknown,
which can be easily solved.

It is the  MAJOR  LESSON  for you from my solution.